shift sum, eve 11/12/99

Patrick Toale (toale@cuhep.Colorado.EDU)
Fri, 12 Nov 1999 22:59:01 -0700 (MST)

Shift Summary for Evening shift, 11/12/99

On Shift: Valeri, Pat

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 8.0
Hours of E799 data (special runs not included): 7.2
Hours of special runs (laser scans, muon runs...): 0.8

Accounting of lost time: TeV problems at begining of shift;
switching to/from muon mode, start/stop runs

========================================================================
Runs ended this shift:
Run 14951 - BEAM_MUON
Run 14952 - BEAM_MUON
Run 14953 - KTEV_PED
Run 14955 - BEAM_799 - 3.9M events

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Description of Shift: pretty smooth data taking with
somewhat low intensity ~ 7.5E12

run spill time comment
------------------------------------------------------------------------
14951 15:30 Shift starts without beam.
15:55 Beam returns. VJ notices we are in 'Access Mode',
call MCR to swich to 'Muon Mode'. End run.

14952 16:05 Start new muon run.
16:15 End muon run.

14953 16:20 Start pedestal run.
16:30 End run. Update pedestals. Some monitors are not
updating, fresh start.

14955 16:40 Start BEAM_799 run 14955.
19:00 Ktevana errors. Instead of kfrend stop run to install
T0's.

14956 19:10 Start run 14956. T0's and SODs are better but
SOD resolution is still bad. Ashkan will install
XTs later after getting more statistics.
20:00 Call MCR about drifting sem, we watch it head
towards 7.0E12. They get it up to around 7.8 at
times but no better.
24:00 Shift ends.

========================================================================
Items to Follow up on Next Shift:
Work on MCR and beam intensity.